# Characterizing Open-Ended Evolution Through Undecidability Mechanisms in Random Boolean Networks

### v1

@@ Abstract
Discrete dynamical models underpin systems biology, but we still lack substrate-agnostic diagnostics for when such models can sustain genuinely open-ended evolution (OEE): the continual production of novel phenotypes rather than eventual settling. We introduce a simple, model-independent metric, $\Omega$, that quantifies OEE as the residence-time-weighted contribution of each attractor’s cycle length across the sequence of attractors realized over time. $\Omega$ is zero for single-attractor dynamics and grows with the number and persistence of distinct cyclic phenotypes, separating enduring innovation from transient noise. Using Random Boolean Networks (RBNs) as a unifying testbed, we compare classical Boolean dynamics with biologically motivated non-classical mechanisms (probabilistic context switching, annealed rule mutation, paraconsistent logic, modal necessary/possible gating, and quantum-inspired superposition/entanglement) under homogeneous and heterogeneous updating schemes. Our results support the view that undecidability-adjacent, state-dependent mechanisms—implemented as contextual switching, conditional necessity/possibility, controlled contradictions, or correlated branching—are enabling conditions for sustained novelty. At the end of our manuscript we outline a practical extension of $\Omega$ to continuous/hybrid state spaces, positioning $\Omega$ as a portable benchmark for OEE in discrete biological modeling and a guide for engineering evolvable synthetic circuits.
@@ Introduction
Biological evolution has generated a broad diversity of forms and functions across scales, from microbes to large multicellular organisms. As an example, if we focus on microbial species alone, their variety has been estimated to be $\sim$$10^{12}$ (locey2016scaling), which is approximately the number of galaxies in the observable universe (conselice2016evolution). One of the central lessons from Darwin1859 is that this diversity is not a static catalogue but is dynamically produced by evolution. Even though any real biosphere is finite and the underlying physical and chemical mechanisms have remained largely unchanged for roughly 4 billion years, the biosphere has continued to explore new combinations of forms and functions without any clear sign of saturation. In this practical sense, biological evolution appears effectively open-ended on geological timescales (packard2019open): the combinatorial space of possible genotypes and phenotypes is so large, and the dynamics of evolution so far from equilibrium, that ongoing novelty can be expected rather than exceptional. In the artificial life literature, this has motivated the study of open-ended evolution as the ongoing production of novel, adaptive structures without an intrinsic upper bound on diversity or complexity (packard2019overview). Since processes having the property of ongoing creative productivity are not necessarily biological, we have generalized this conceptualization to include open-endedness in any form (stanley2019open).
@@
Despite its ubiquity in the literature, there is no full consensus on how to define open-ended evolution (OEE). Some authors take the continual generation of novelty—the sustained appearance of previously unseen entities or behaviors—to be sufficient for a system to be labeled as open-ended (soros2014identifying). Others argue that OEE should be reserved for cases in which some notion of system complexity can grow without an a priori upper bound, as seems to be suggested by empirical scaling laws across levels of biological organization (corominas2018zipf). However, this immediately raises the methodological problem of how to define and measure complexity in a way that is both domain-independent and monotonically increasing along evolutionary trajectories. As a result, the literature distinguishes between variety-based and complexity-based conceptions, and more generally between multiple “flavors” of open-endedness (taylor2016open). For our own purposes, we will operationally define open-endedness as the capability of a system to continually generate novel (without necessarily being increasingly complex) structures or behaviors without settling into fixed or cyclic states. In this way, given a system, we are interested in how to classify its dynamics as open-ended.
@@
So far, two overarching suites of metrics have been proposed to measure OEE. On the one hand, Evolutionary Activity Statistics measures the degree of adaptive dynamics within a population by tracking the persistence of specific elements, called “components” (bedau1997comparison; bedau1998classification). As a disadvantage of this approach, we should know what the “components” are ab initio; these, in fact, are selected arbitrarily. Furthermore, it also fails to take into account that “components” can evolve over time, disregarding their ability to acquire new functions (or become obsolete) through adaptation (wong2023roles). On the other hand, Measurements of Open-ended Dynamics in Evolving Systems (MODES) is a toolbox designed to analyze and quantify the dynamics of evolving systems, specifically focusing on identifying features that are indicative of OEE (dolson2019modes). Although MODES is an ambitious proposal, its application is limited to specific systems. In fact, some of its metrics only make sense in AVIDA, an artificial life software platform to study the evolutionary biology of self-replicating and evolving digital organisms (lenski2003evolutionary).
@@
Moreover, classifying a system as open-ended does not help us to understand how this property arose evolutionary speaking. As stepney2024open have pointed out, the focus should be on scrutinizing the mechanisms underlying open-endedness, not simply trying to quantify it. Of course this requires a thinking shift with respect the current approaches to OEE, suggesting that the capacity for open-endedness arises from evolution itself, rather than being a pre-existing condition of the entire evolutionary system. This perspective, championed by pattee2019evolved, emphasizes how organisms can create their own mechanisms, such as language, that make late evolution more open-ended. Thus, we must first discern what universal conditions or mechanisms contribute to the emergence of evolutionary open-endedness.
@@
Using a sufficiently general and formal approach to circumscribe any resource-bounded discrete system, adams2017formal have shown that state-dependence is a universal mechanism to achieve OEE. However, as the authors recognized, they did not explore self-referential mechanisms between organisms and their environment, which they warned could have enriched their experimental findings. As we have shown in previous research using the power and generality of category theory to encompass all kinds of mathematical structures (and formal or computational models), semantic closure—the self-referential mechanism where an organism’s own actions and interpretations create the context for future actions—is crucial for robust self-replication, niche construction, and evolved open-endedness in any system (10.1098/rsif.2025.0784).
@@
In particular, in the case of non-cyclical discrete dynamical systems where there are an infinite number of possible states and each state must contain all the information needed to compute successive states, hernandez2018undecidability have formally proved undecidability to be a requirement for OEE, showing that adaptation itself is a generalization of the halting problem (turing1936computable). The existence of adapted behaviors whose appearance is formally undecidable reframes evolution as not merely a search through a possibility space, but as an ongoing construction of uncomputable futures (kauffman2021world). Therefore, any mechanism associated to the impossibility of constructing a single algorithm that always leads to a correct yes-or-no answer is a potential candidate for promoting open-endedness.
@@
Using Random Boolean Networks (RBNs), we build on earlier work that treated changes in the attractor landscape as a proxy for evolvability. In particular, TorresSosa2012 introduced selection criteria that reward networks which both conserve existing attractors and generate new ones, showing that this conservation-innovation trade-off can drive the evolution of critical dynamics. Here we repurpose this attractor-level perspective as an explicit, substrate agnostic metric of open-endedness that is applicable to any dynamical system with discrete state space. Recognizing that open-endedness is inherently a product of evolutionary processes, we further explore the underlying mechanisms driving this phenomenon. Particularly, we focus our attention on non-classical logics: formal systems that modify, extend, or deviate from the foundational principles of classical logic by rejecting or altering concepts such as the law of excluded middle (a statement is either true or false) or bivalence (only two truth values exist) (priest2008introduction). These alternative frameworks are necessary for modeling situations where classical logic is insufficient, such as dealing with incomplete information, statements that are neither true nor false, paradoxical information, and alternative kinds of knowledge and reasoning. In this way, we investigate the role of undecidability in its different facets to promote open-endedness in a biological model.
@@
In the next section, we detail our methodology, emphasizing the relevance of Random Boolean Networks (RBNs) and non-classical logics in biological modeling, while also introducing our novel open-endedness metric. Afterwards, we proceed to show our results considering both homogeneous and heterogeneous RBNs, exploring whether structural and temporal heterogeneity also serve as key drivers of open-endedness. Subsequently, we discuss the results obtained, revealing strategic insights into how systems can maximize their open-endedness based on their inherent conditions. Finally, we conclude our work by proposing avenues for future research, including a generalization of our metric to encompass systems with continuous state space.
@@ Methods
A gene regulatory network (GRN) is a collection of genes, RNA, and proteins that control which genes are turned on or off, and how active they are, in a cell (davidson2005gene). They act as a cellular control system, determining when and where genes are expressed, which in turn dictates the cell’s function and behavior. These networks are crucial for fundamental biological processes like cell development (levine2005gene), differentiation (roy2014gene), and responding to environmental changes (crombach2008evolution). As a way of capturing the complexity of GRNs, Kauffman1969 first introduced Random Boolean Networks (RBNs): discrete dynamical systems modeled as directed networks, where each node (a gene) has a Boolean value (0 or 1) and is randomly connected to other nodes.
Formally, a directed graph $D$ consists of a set $V$ of elements $a,b,c,\dots$ called the nodes of $D$ and a set $A$ of ordered pairs of nodes $(a,b)$, $(b,c)$, $\dots$ called the arcs of $D$. We use the symbol $ab$ to represent the arc $(a,b)$. If $ab$ is in the arc set $A$ of $D$, then we say that $a$ is an in-neighbor of $b$, and also that $b$ is an out-neighbor of $a$. We say that $D$ is $k$-in regular $(k\geq 1)$ if every node has exactly $k$ in-neighbors. The out-degree of a node $a$ is the number of out-neighbors of $a$. Let $D=(V,A)$ be a $k$-in directed graph; a family $\left(f_{a}\right)_{a\in V}$ of functions $f_{a}:\{0,1\}^{k}\longrightarrow\{0,1\}$ is called a Boolean network on $D$.
A Boolean network is called random if the assignment $a\mapsto f_{a}$ is made at random by sampling independently and uniformly from the set of all the $2^{2^{K}}$ Boolean functions with $K$ inputs. A function $\phi:V\longrightarrow\{0,1\}$, $a\mapsto\phi_{a}$, is called a state of the random Boolean network on $D$. The value $\phi_{a}$ is called the state of $a$. The updating function $F(\phi)$ of a state $\phi$ is the function $F:V\longrightarrow\{0,1\}$, $a\mapsto\phi^{\prime}_{a}$, defined as
$\phi^{\prime}_{a}=f_{a}\left(\phi_{a_{1}},\dots,\phi_{a_{k}}\right).$ (1)
Given the complexity exhibited by real GRNs, Kauffman assumed the following four conditions in his model (Kauffman1969):
  1. 1.
    Each gene is regulated by exactly $K$ other genes.
  2. 2.
    The $K$ genes that regulate each node are chosen randomly using a uniform probability.
  3. 3.
    All genes are updated synchronously, i.e., at the same time step  (Gershenson2004b).
  4. 4.
    Each gene $f_{a}$ maps the input $\left(\phi_{a_{1}},\dots,\phi_{a_{k}}\right)$ to the ‘on’ state with probability $p$ and to the ‘off’ state with probability 1p1-p.
The first two imply structural homogeneity, the third temporal homogeneity, and the fourth functional homogeneity. In this paper, a heterogeneous RBN refers to an architecture with structural and temporal heterogeneity (sanchez2023heterogeneity). Otherwise, a homogeneous RBN refers to an architecture with structural, temporal, and functional homogeneity. Following the methodology proposed by lopez2023temporal, to implement structural homogeneity (heterogeneity), we use Poisson (Exponential) as the degree distribution. For implementing temporal heterogeneity we resort to the pre-programmed asynchronous scheme in CUBEWALKERS, an open-source, GPU-based Python tool for simulating Boolean networks (park2023models). Once we have introduced RBNs, let us discuss the biological role of non-classical logics and how we can implement them using RBNs.
@@
A PBN is a collection of Boolean networks in which a constituent network governs gene activity for a random period of time before another randomly chosen constituent network takes over. Stochastic phenotype switching is a well-established adaptation strategy in fluctuating environments, classically demonstrated in many bacterial systems (acar2008stochastic). PBNs were introduced precisely to capture rule-based uncertainty and context switching in GRNs, and have since become a standard for modeling, inference, and control of genetic networks (shmulevich2010probabilistic). This makes PBNs the most direct way to incorporate environmental contingency and regulatory program switching into an otherwise Boolean framework.
Formally, a PBN consists of a sequence $V=\{x_{i}\}_{i=1}^{n}$ of $n$ nodes, where $x_{i}\in\{0,1\}$, and a sequence $\{\textbf{f}_{l}\}_{l=1}^{m}$ of vector-valued functions, defining constituent networks. Each vector-valued function $\textbf{f}_{l}=(f_{l}^{(1)},f_{l}^{(2)},...,f_{l}^{(n)})$ determines a constituent network, or context, of the PBN. The function $f_{l}^{(i)}:\{0,1\}^{k}\rightarrow\{0,1\}$ is a predictor of gene $i$ whenever network $l$ is selected. At each updating epoch a decision is made whether to switch the constituent network. A constituent network is randomly selected from among all constituent networks according to the selection probability distribution $\{c_{l}\}_{l=1}^{m}$ such that $\sum_{l=1}^{m}c_{l}=1$. In our code, we generate $m$ pure-Boolean contexts by resampling lookup tables (LUTs) from a base BN and set uniform $c_{l}$. We then execute long segments in the current context, updating deterministically within that context and only sampling a new context at segment boundaries. This cleanly reproduces the standard PBN semantics used in the gene-regulation literature (shmulevich2010probabilistic).
@@
GRN logic is not immutable: transcriptional programs can rewire or tune their effective rules across time due to chromatin remodeling (filtz2014regulation), TF post-translational modifications (swift2017matter), or changing cofactor abundance (tomljanovic2025epigenetic). From a modeling standpoint, this is the annealed perspective familiar from RBN theory, where rule tables are treated as stochastic variables across updates, capturing plasticity not represented by fixed (“quenched”) rules. Empirically, modern time-resolved genomics underscores that regulatory interactions are dynamic over minutes-hours, supporting annealed or partially annealed abstractions when studying innovation and adaptability (meeussen2024time). The annealed approximation of derrida1986random remains a standard conceptual scaffold, and recent reviews emphasize explicitly time-dependent GRN modeling (marku2023time).
@@
Cells frequently face conflicting regulatory cues: the same upstream factor can both promote and inhibit a response depending on context, dosage, or timing (hart2012design). This is well documented in immune and developmental systems, e.g., IL-2 simultaneously supporting proliferation and death in T cells (hart2013utility); such circuits exploit antagonistic arms to achieve robustness, fold-change detection, and pulsatile dynamics, hence called “paradoxical components” (hart2014paradoxical). These phenomena motivate logics that tolerate local contradictions rather than exploding to triviality, which is exactly what paraconsistent formalisms provide (Gershenson1999). Using an inconsistency-tolerant semantics at the rule/local-gate level mirrors how GRNs integrate pleiotropic, noisy, or partially contradictory inputs without collapsing function.
@@
Many regulatory claims in biology are modal: a factor may be necessary (must be present) or merely possible/sufficient (can enable) for an outcome, and these claims are inherently contextual (cell type, chromatin state, signaling milieu) (monteiro2008temporal). The systems biology community routinely encodes such “must/possibly” properties with temporal/modal logics and verifies them on GRN models via model checking, e.g., verifying that “eventually gene A activates B and always C stays off” in stress-response or cell-cycle networks (monteiro2014model). This makes modal operators a natural abstraction layer for contextual constraints in GRN dynamics. Representative overviews and applications include model checking for logical regulatory networks and temporal-logic pattern mining in cellular models (giacobbe2015model).
@@
Formally, a modalized RBN augments a standard RBN with a Kripke frame $\mathcal{K}=(V,R)$, where $V$ is the set of nodes and $R\subseteq V\times V$ encondes accessibility among nodes. Each LUT entry for node $i$ may take values in $\{0,1,\text{possible},(\text{necessary},v)\}$. At time $t$, let $A_{i}(t)=\{j\in V|(i,j)\in R\}$ be $i$’s accesible set and let $\textbf{x}_{A_{i}}(t)$ be their current states. The evaluation map $\mu$ is: $\mu(\text{possible})=1$ iff $\exists j\in A_{i}(t):x_{j}(t)=1$; $\mu\left((\text{necessary},v)\right)=v$ iff $\forall j\in A_{i}(t):x_{j}(t)=1$, else 0; and $\mu(b)=b$ for $b\in\{0,1\}$. The network update is $x_{i}(t+1)=\mu(f_{i}(x_{i}(t)))$. In our code, we build a Kripke frame as described above and sample an accessibility degree $d$, wiring each node to up to $d$ accessible nodes. It then rewrites LUT outputs by flipping entries to the literal tokens possible or necessary (the latter carrying a concrete bit) with probabilities $p_{\text{possible}}$ and $p_{\text{necessary}}$. This cleanly imports necessity/possibility semantics into our RBN while keeping the base state space Boolean.
Quantum Logic

### v2

@@ Abstract
Discrete dynamical models underpin systems biology, but we still lack substrate-agnostic diagnostics for when such models can sustain genuinely open-ended evolution (OEE): the continual production of novel phenotypes rather than eventual settling. We introduce a simple, model-independent metric, $\Omega$, that quantifies OEE as the residence-time-weighted average of attractor cycle lengths across the sequence of attractors realized over time. $\Omega$ is zero for single-attractor dynamics and grows with the number and persistence of distinct cyclic phenotypes, separating enduring innovation from transient noise. Using Random Boolean Networks (RBNs) as a unifying testbed, we compare classical Boolean dynamics with biologically motivated non-classical mechanisms (probabilistic context switching, annealed rule mutation, paraconsistent logic, modal necessary/possible gating, and quantum-inspired superposition/paired-state coupling) under homogeneous and heterogeneous updating schemes. Our results support the view that undecidability-adjacent, state-dependent mechanisms—implemented as probabilistic context switching, modal necessity/possibility gating, paraconsistent logic (controlled contradictions), or quantum-inspired superposition/paired-state coupling (correlated branching)—are enabling conditions for sustained novelty. At the end of our manuscript we outline a practical extension of $\Omega$ to continuous/hybrid state spaces, positioning $\Omega$ as a portable benchmark for OEE in discrete biological modeling and a guide for engineering evolvable synthetic circuits.
@@ Introduction
Biological evolution has generated a broad diversity of forms and functions across scales, from microbes to large multicellular organisms. As an example, if we focus on microbial species alone, their variety has been estimated to be $\sim$$10^{12}$ (Locey and Lennon, 2016), which is approximately the number of galaxies in the observable universe (Conselice et al., 2016). One of the central lessons from Darwin (1859) is that this diversity is not a static catalogue but is dynamically produced by evolution. Even though any real biosphere is finite and the underlying physical and chemical mechanisms have remained largely unchanged for roughly 4 billion years, the biosphere has continued to explore new combinations of forms and functions without any clear sign of saturation. In this practical sense, biological evolution appears effectively open-ended on geological timescales (Packard et al., 2019b): the combinatorial space of possible genotypes and phenotypes is so large, and the dynamics of evolution so far from equilibrium, that ongoing novelty can be expected rather than exceptional. In the artificial life literature, this has motivated the study of open-ended evolution as the ongoing production of novel, adaptive structures without an intrinsic upper bound on diversity or complexity (Packard et al., 2019a). Since processes having the property of ongoing creative productivity are not necessarily biological, we have generalized this conceptualization to include open-endedness in any form (Stanley, 2019).
@@
Despite its ubiquity in the literature, there is no full consensus on how to define open-ended evolution (OEE). Some authors take the continual generation of novelty—the sustained appearance of previously unseen entities or behaviors—to be sufficient for a system to be labeled as open-ended (Soros and Stanley, 2014). Others argue that OEE should be reserved for cases in which some notion of system complexity can grow without an a priori upper bound, as seems to be suggested by empirical scaling laws across levels of biological organization (Corominas-Murtra et al., 2018). However, this immediately raises the methodological problem of how to define and measure complexity in a way that is both domain-independent and monotonically increasing along evolutionary trajectories. As a result, the literature distinguishes between variety-based and complexity-based conceptions, and more generally between multiple “flavors” of open-endedness (Taylor et al., 2016). For our own purposes, we will operationally define open-endedness as the capability of a system to continually generate novel (without necessarily being increasingly complex) structures or behaviors without settling into fixed or cyclic states. In this way, given a system, we are interested in how to classify its dynamics as open-ended.
@@
So far, two overarching suites of metrics have been proposed to measure OEE. On the one hand, Evolutionary Activity Statistics (EAS) measure the degree of adaptive dynamics within a population by tracking the persistence of specific elements, called “components” (Bedau et al., 1997, 1998). As a disadvantage of this approach, we should know what the “components” are ab initio; these, in fact, are selected arbitrarily. Furthermore, it also fails to take into account that “components” can evolve over time, disregarding their ability to acquire new functions (or become obsolete) through adaptation (Wong et al., 2023). On the other hand, Measurements of Open-ended Dynamics in Evolving Systems (MODES) is a toolbox designed to analyze and quantify the dynamics of evolving systems, specifically focusing on identifying features that are indicative of OEE (Dolson et al., 2019). Although MODES is an ambitious proposal, its application is limited to specific systems. In fact, some of its metrics only make sense in AVIDA, an artificial life software platform to study the evolutionary biology of self-replicating and evolving digital organisms (Lenski et al., 2003).
@@
Moreover, classifying a system as open-ended does not help us to understand how this property arose evolutionary speaking. As Stepney and Hickinbotham (2024) have pointed out, the focus should be on scrutinizing the mechanisms underlying open-endedness, not simply trying to quantify it. Of course this requires a thinking shift with respect the current approaches to OEE, suggesting that the capacity for open-endedness arises from evolution itself, rather than being a pre-existing condition of the entire evolutionary system. This perspective, championed by Pattee and Sayama (2019), emphasizes how organisms can create their own mechanisms, such as language, that make late evolution more open-ended. Thus, we must first discern what universal conditions or mechanisms contribute to the emergence of evolutionary open-endedness.
@@
Using a sufficiently general and formal approach to circumscribe any resource-bounded discrete system, Adams et al. (2017) have shown that state-dependence is a universal mechanism to achieve OEE. However, as the authors recognized, they did not explore self-referential mechanisms between organisms and their environment, which they warned could have enriched their experimental findings (Adams et al., 2017, p. 13-14). As we have shown in previous research using the power and generality of category theory to encompass all kinds of mathematical structures (and formal or computational models), semantic closure—the self-referential mechanism where an organism’s own actions and interpretations create the context for future actions—is crucial for robust self-replication, niche construction, and evolved open-endedness in any system (López-Díaz and Gershenson, 2025).
@@
In particular, in the case of non-cyclical discrete dynamical systems where there are an infinite number of possible states and each state must contain all the information needed to compute successive states, Hernández-Orozco et al. (2018) have formally proved undecidability to be a requirement for OEE, showing that adaptation itself is a generalization of the halting problem (Turing, 1936). The existence of adapted behaviors whose appearance is formally undecidable reframes evolution as not merely a search through a possibility space, but as an ongoing construction of uncomputable futures (Kauffman and Roli, 2021). Therefore, any mechanism associated to the impossibility of constructing a single algorithm that always leads to a correct yes-or-no answer is a potential candidate for promoting open-endedness.
@@
Using Random Boolean Networks (RBNs), we build on earlier work that treated changes in the attractor landscape as a proxy for evolvability. In particular, Torres-Sosa et al. (2012) introduced selection criteria that reward networks which both conserve existing attractors and generate new ones, showing that this conservation-innovation trade-off can drive the evolution of critical dynamics. Here we repurpose this attractor-level perspective as an explicit, substrate agnostic metric of open-endedness that is applicable to any dynamical system with discrete state space. Recognizing that open-endedness is inherently a product of evolutionary processes, we further explore the underlying mechanisms driving this phenomenon. Particularly, we focus our attention on non-classical logics: formal systems that modify, extend, or deviate from the foundational principles of classical logic by rejecting or altering concepts such as the law of excluded middle (a statement is either true or false) or bivalence (only two truth values exist) (Priest, 2008). These alternative frameworks are necessary for modeling situations where classical logic is insufficient, such as dealing with incomplete information, statements that are neither true nor false, paradoxical information, and alternative kinds of knowledge and reasoning. In this way, we investigate the role of undecidability in its different facets to promote open-endedness in a biological model.
From a systems-biology standpoint, the practical value of $\Omega$ is that it can be computed for any existing discrete state space model, not only for random ensembles. Given a curated Boolean regulatory circuit, one can simulate long trajectories under experimentally motivated sources of variability (e.g., asynchronous update timing, fluctuating inputs, or context switches) and compute $\Omega$ over successive windows. A high $\Omega$ indicates that the model repeatedly visits multiple long-lived expression programs, rather than merely flickering among transient states or collapsing to a single fate. This offers a compact, model-agnostic summary of phenotypic plasticity that can be used to compare alternative network reconstructions, to quantify how mutations or interventions reshape the accessible repertoire of stable phenotypes, and to guide the design of synthetic circuits intended to support rich but structured switching.
This work has two distinct aims that we keep conceptually separate. First, we define and motivate $\Omega$ as a finite-horizon, recurrence-based diagnostic of sustained novelty in discrete dynamical systems, and we specify exactly how it is computed from empirical recurrence episodes (including stochastic dynamics). Second, we use $\Omega$ as an analysis lens for comparing classes of logic-inspired update mechanisms (context switching, conditional gating, controlled contradictions, and correlated branching) within a common Random Boolean Network testbed.
To make this separation explicit, our validation proceeds in four steps. (i) We establish analytic sanity checks showing that $\Omega$ vanishes for single-attractor deterministic dynamics and also for “pure noise” dynamics with bounded dwell times (see Methods). (ii) We provide an operational definition of recurrence episodes (cycle discovery, dwell time, escape) that makes clear how chained cycles and non-Boolean intermediate tokens are treated. (iii) We quantify horizon sensitivity and size effects by computing $\Omega$ across multiple observation windows and including a small-$N$ control (Supplementary Fig. S22). (iv) We validate portability beyond random RBN ensembles by evaluating $\Omega$ on external benchmarks (Elementary Cellular Automata and curated Boolean GRN models), and we compare $\Omega$ to standard trajectory-level baselines (unique-state fraction, node entropy, and a compression proxy) together with a complement-symmetry control (Supplementary Figs. S23–S25). Together these checks support interpreting $\Omega$ as a finite-time signature of recurrence-weighted sustained novelty, rather than as a strict $T\to\infty$ attractor decomposition for stochastic dynamics.
@@
In the next section, we detail our methodology, emphasizing the relevance of Random Boolean Networks (RBNs) and non-classical logics in biological modeling, while also introducing our open-endedness metric. Afterwards, we proceed to show our results considering both homogeneous and heterogeneous RBNs, exploring whether structural and temporal heterogeneity also serves as key drivers of open-endedness. Subsequently, we discuss the results obtained, revealing strategic insights into how systems can maximize their open-endedness based on their inherent conditions and properties. Finally, we conclude our work by proposing avenues for future research, including a generalization of our metric to encompass systems with continuous state space.
@@ Methods
Boolean network models provide a widely used coarse-grained description of gene regulatory networks (GRNs) in which each gene is represented as an “active”/“inactive” variable and regulation is encoded as a directed interaction graph (Kauffman, 1969; Gershenson, 2004a). We use Random Boolean Networks (RBNs) as a controlled ensemble in which wiring, logic, and update timing can be systematically manipulated while retaining the standard biological interpretation of attractors as putative cell fates.
An RBN is specified by $N$ nodes, an in-neighborhood $I_{i}$ for each node $i$ (with $|I_{i}|=k_{i}$), and a Boolean update rule $f_{i}:\{0,1\}^{k_{i}}\to\{0,1\}$ (conveniently represented as a lookup table, alias LUT). Writing the network state as $\mathbf{x}(t)=(x_{1}(t),\ldots,x_{N}(t))\in\{0,1\}^{N}$, synchronous dynamics are given by
$x_{i}(t+1)=f_{i}\!\left(\mathbf{x}_{I_{i}}(t)\right).$ (1)
In the classical Kauffman ensemble, regulators and truth tables are sampled at random; in our simulations we draw regulators according to a chosen in-degree distribution and sample each lookup table with bias $p$ (we use $p=0.5$ unless stated otherwise).
To probe architectural effects, we consider two regimes. In the homogeneous regime we use a narrow in-degree distribution (Poisson with mean $K$), synchronous updates, and fixed (quenched) truth tables. In the heterogeneous regime we introduce structural and temporal heterogeneity by using a broader in-degree distribution (Exponential with the same mean $K$) together with stochastic asynchronous updates implemented with CUBEWALKERS (Park et al., 2023). This heterogeneity protocol follows López-Díaz et al. (2023); here we summarize only what is needed to reproduce our simulations and focus on how different state-dependent mechanisms modify the dynamics. Once we have introduced RBNs, let us discuss the biological role of non-classical logics and how we can implement them using RBNs.
@@
Stochastic phenotype switching is a well-established adaptation strategy in fluctuating environments, classically demonstrated in many bacterial systems (Acar et al., 2008). PBNs were introduced precisely to capture rule-based uncertainty and context switching in GRNs, and have since become a standard for modeling, inference, and control of genetic networks (Shmulevich and Dougherty, 2010). A PBN is a collection of Boolean networks in which a constituent network governs gene activity for a random period of time before another randomly chosen constituent network takes over. This makes PBNs the most direct way to incorporate environmental contingency and regulatory program switching into an otherwise Boolean framework.
Formally, a PBN consists of a sequence $V=\{x_{i}\}_{i=1}^{n}$ of $n$ nodes, where $x_{i}\in\{0,1\}$, and a sequence $\{\textbf{f}_{l}\}_{l=1}^{m}$ of vector-valued functions, defining constituent networks. Each vector-valued function $\textbf{f}_{l}=(f_{l}^{(1)},f_{l}^{(2)},...,f_{l}^{(n)})$ determines a constituent network, or context, of the PBN. The function $f_{l}^{(i)}:\{0,1\}^{k}\rightarrow\{0,1\}$ is a predictor of gene $i$ whenever network $l$ is selected. At each updating epoch a decision is made whether to switch the constituent network. A constituent network is randomly selected from among all constituent networks according to the selection probability distribution $\{c_{l}\}_{l=1}^{m}$ such that $\sum_{l=1}^{m}c_{l}=1$. In our code, we generate $m$ pure-Boolean contexts by resampling lookup tables (LUTs) from a base Boolean Network (BN) and set uniform $c_{l}$. We then execute long segments in the current context, updating deterministically within that context and only sampling a new context at segment boundaries. This cleanly reproduces the standard PBN semantics used in the gene-regulation literature (Shmulevich and Dougherty, 2010).
@@
GRN logic is not immutable: transcriptional programs can rewire or tune their effective rules across time due to chromatin remodeling (Filtz et al., 2014), transcription factor post-translational modifications (Swift and Coruzzi, 2017), or changing cofactor abundance (Tomljanović et al., 2025). From a modeling standpoint, this is the annealed perspective familiar from RBN theory, where rule tables are treated as stochastic variables across updates, capturing plasticity not represented by fixed (“quenched”) rules. Empirically, modern time-resolved genomics underscores that regulatory interactions are dynamic over minutes-hours, supporting annealed or partially annealed abstractions when studying innovation and adaptability (Meeussen and Lenstra, 2024). The annealed approximation of Derrida and Pomeau (1986) remains a standard conceptual scaffold, and recent reviews emphasize explicitly time-dependent GRN modeling (Marku and Pancaldi, 2023).
@@
Cells frequently face conflicting regulatory cues: the same upstream factor can both promote and inhibit a response depending on context, dosage, or timing (Hart et al., 2012). This is well documented in immune and developmental systems, e.g., IL-2 simultaneously supporting proliferation and death in T cells (Hart and Alon, 2013); such circuits exploit antagonistic arms to achieve robustness, fold-change detection, and pulsatile dynamics, hence called “paradoxical components” (Hart et al., 2014). These phenomena motivate logics that tolerate local contradictions rather than exploding to triviality, which is exactly what paraconsistent formalisms provide (Gershenson, 1999). Using an inconsistency-tolerant semantics at the rule/local-gate level mirrors how GRNs integrate pleiotropic, noisy, or partially contradictory inputs without collapsing function.
@@
Many regulatory claims in biology are modal: a factor may be necessary (must be present) or merely possible/sufficient (can enable) for an outcome, and these claims are inherently contextual (cell type, chromatin state, signaling milieu) (Monteiro et al., 2008). The systems biology community routinely encodes such “must/possibly” properties with temporal/modal logics and verifies them on GRN models via model checking, e.g., verifying that “eventually gene A activates B and always C stays off” in stress-response or cell-cycle networks (Monteiro et al., 2014). This makes modal operators a natural abstraction layer for contextual constraints in GRN dynamics. Representative overviews and applications include model checking for logical regulatory networks and temporal-logic pattern mining in cellular models (Giacobbe et al., 2015).
@@
Formally, a modalized RBN augments a standard RBN with a Kripke frame $\mathcal{K}=(V,R)$, where $V$ is the set of nodes and $R\subseteq V\times V$ encodes accessibility among nodes. Each LUT entry for node $i$ may take values in $\{0,1,\text{possible},(\text{necessary},v)\}$. At time $t$, let $A_{i}(t)=\{j\in V|(i,j)\in R\}$ be $i$’s accessible set, and let $\textbf{x}_{A_{i}}(t)$ be their current states. The evaluation map $\mu$ is: $\mu(\text{possible})=1$ iff $\exists j\in A_{i}(t):x_{j}(t)=1$; $\mu\left((\text{necessary},v)\right)=v$ iff $\forall j\in A_{i}(t):x_{j}(t)=1$, else 0; and $\mu(b)=b$ for $b\in\{0,1\}$. The network update is $x_{i}(t+1)=\mu(f_{i}(x_{i}(t)))$. In our code, we build a Kripke frame as described above and sample an accessibility degree $d$, wiring each node to up to $d$ accessible nodes. It then rewrites LUT outputs by flipping entries to the literal tokens possible or necessary (the latter carrying a concrete bit) with probabilities $p_{\text{possible}}$ and $p_{\text{necessary}}$. This cleanly imports necessity/possibility semantics into our RBN while keeping the base state space Boolean.
Quantum-inspired correlated branching